OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

OpenStudy (nincompoop):

FIRST STEP IS TO ASSIGN WHICH IS YOUR POINT A AND WHICH IS POINT B LABEL YOUR COORDINATES A(X1, Y1); B(X2, Y2) OBTAIN THE SLOPE OF THE LINE SLOPE = (Y2-Y1)/(X2-X1) THEN USE THE POINT-SLOPE INTERCEPT FORM Y-Y1 = M(X-X1) WHERE M IS YOUR SLOPE USE DISTRIBUTIVE PROPERTY FOR THE RIGHT-HAND SIDE, THEN SOLVE FOR Y

OpenStudy (afifarahman):

as slope, m = -1 so , \[y=-x+9\]

OpenStudy (nincompoop):

AFIFICOOCOO, PLEASE DO NOT GIVE OUT ACTUAL ANSWERS IN THE FUTURE. CHEATING IS NOT GOOD, COOCOOBIRD